OpenStudy (anonymous):

how do you solve The seats at a local baseball stadium are arranged so that each row has five more seats than the row in front of it. If there are four seats in the first row, how many total seats are in the first 24 rows?

OpenStudy (anonymous):

You have an arithmetic series a1=4, a2=4+5=9= a1+5 a3=9+5=14=a1+2(5) ................. an=a1+(n-1)(5) n=24 a_(24)=a1+23*5=4+23*5=119 Sn= n(a1+an)/ 2= S_(24)= 24(4+119)/2=1476

OpenStudy (anonymous):

okay, what does that mean

OpenStudy (anonymous):

Guykubpommm helpppp me!!

OpenStudy (anonymous):

a1 = 4 and a2 = 4 +5,a3 = 4 +2*5 and then an = 4+(n-1)*5 so the sum is sigma(4+(n-1)*5) when n runs from 1 to 24

OpenStudy (anonymous):

sigma(4+(n-1)*5) = sigma (4n-1)

OpenStudy (anonymous):

so it is sigma4n - sigma1 = 4*24*25/2 - 24

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yeah!! is it 1176

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

YESSSSS. Thankkk you! your ammazing haha

OpenStudy (anonymous):

thanks lol

OpenStudy (anonymous):

no worries! haha

OpenStudy (anonymous):

but no I figured it wrong haha

OpenStudy (anonymous):

it is sigma(5n-1) = 5*24*25/2 - 24 = 1476 lol

OpenStudy (anonymous):

yeahh I know but I still got it right on my work cause I just guessed the closet answer to the answer we figured the first time lol sooo its okay